OpenStudy (anonymous):

The marginal revenue (in thousands of dollars) from the sale of x tractors is R'(x)=2x(x^2+10)^2/1000 Find the revenue function R if the revenue from the sale of 4 tractors is $198,000.

OpenStudy (dumbcow):

ooh a economics question...kinda basically you have to integrate R'(x) to find R(x) do this by substituting u = x^2 + 10 du = 2x dx --> dx = du/2x leaving \[\frac{1}{1000}\int\limits_{}^{}u^{2} du\]